OpenStudy (anonymous):

3n^2+sqrt(2n^2)/sqrt(10n) I'm working on simplifying radical expressions. I can do the more simple problems but this one is stumping me a little. I know the answer but would like to see how you got there. Thanks so much!

OpenStudy (anonymous):

\[3n ^{2}+\sqrt{2n ^{2}}/\sqrt{10n}\]

OpenStudy (anonymous):

\[3n ^{2}+\sqrt{n/5}\]